OpenStudy (anonymous):

i am compleately stumped on this one.. pleaase help Find the exact area of a circle having the given circumference. 4(sqareroot3pi) A =

OpenStudy (anonymous):

\(c=\sqrt3\times\pi\) or \(c = \sqrt{3\pi}\)

OpenStudy (anonymous):

\(c = d \pi = 2\pi r\)

OpenStudy (anonymous):

the answer are 2sqrt3pi 4sqrt3pi or 12pi

OpenStudy (anonymous):

I'm still not sure of the question. Is \(\pi\) under the radical or not?

OpenStudy (anonymous):

Use the circumference formula to find the radius, then use the radius to find the area.